I fitzed with the PRM and SRM briefly, and I now believe that they're both ready to go into the chambers.
For each optic, I used the microscope on a micrometer to check that the scribe lines on each side of the optic were at the same height. Basic procedure was to center the microscope on one scribe line, move the microscope to the other side, to see how far the line was from center, and try to (very gently!!) rotate the optic in the wire about the z-axis about half the distance that the one scribe line needed to be. Rinse and repeat several times until satisfied.
I then checked that our HeNe oplev was still at 5.5" beam height, and that the beam traveled straight across the table. I put the SRM in the oplev, unclamped the EQ stops, and waited for it to settle. The HEPA filters were turned off, to minimize the breeze. While the SRM settled, I worked on the height/rotation for the PRM on the other table.
After checking the SRM balance, I clamped it and moved it, and checked the PRM balance, then turned off the HeNe and rewrapped everything in foil, and turned on the HEPAs.
Both the SRM and the PRM seem a little off in Pitch. The beam returning to the QPD (placed just next to the laser) was always ~1cm above the center of the QPD. The beam travel distance was ~3m (vaguely) from laser to optic to QPD. This effect may be because the optics were originally balanced with OSEMs in place, and I didn't have any OSEMs today. Koji and I found several months ago that the OSEMs have some DC affect on the optics.
Anyhow, since our optics are so small, I think the OSEMs and coils can handle this small DC offset in pitch, so I think we're ready to rock-n-roll with putting them in the chambers.
Still on the to-do list......Tip Tilts!
The photo shows the oplev beam position on (kind of) the QPD, for the SRM. The PRM was basically the same.